Could you please help me with this problem.
Write the equation of an elipse given vertex (15,0) and focus(12,0)

Answer by mukhopadhyay(490)

There are two vertices of an ellipse, and they are equidistant from the center.
The center of the ellipse having vertices at (-15,0) and (15,0) is (0,0). The major axis is x-axis.
Standard equation is:
x^2/a^2 + y^2/b^2 = 1
a=15 (distance between center and vertex)
c=12 (distance between center and focus)
c^2=a^2-b^2
=>b^2=15^2-12^2=27*3=81
=>b=9
So, the equation of the ellipse is:
x^2/225+y^2/81=1